# Leveraging Pre-Trained Visual Transformers for Multi-Band Photometric Light Curve Classification

### v2

@@ Abstract
Aims. This study investigates the potential of a pre-trained visual transformer (VT) model, specifically the Swin Transformer V2 (SwinV2), to classify photometric light curves without the need for feature extraction or multi-band preprocessing. The goal is to assess whether this image-based approach can accurately differentiate astronomical phenomena and if it can serve as a viable option for working with multi-band photometric light curves.
Methods. We transformed each multi-band light curve into an image. These images serve as input to the SwinV2 model, which is pre-trained on ImageNet-21K. The datasets employed include the public Catalog of Variable Stars from the Massive Compact Halo Object (MACHO) survey, using both one and two bands, and the first round of the recent Extended LSST Astronomical Time-Series Classification Challenge (ELAsTiCC), which includes six bands. The model’s performance was evaluated on 6 classes for the MACHO dataset and 20 distinct classes of variable stars and transient events for the ELAsTiCC dataset.
Results. The fine-tuned SwinV2 achieved better performance than models specifically designed for light curves, such as Astromer and the Astronomical Transformer for time series And Tabular data (ATAT). When trained on the Full MACHO dataset, it attained a macro F1-score of 80.2% and outperformed Astromer in single-band experiments. Incorporating a second band further improved performance, increasing the F1-score to $84.1\%$. In the ELAsTiCC dataset, SwinV2 achieved a macro F1-score of $65.5\%$, slightly surpassing ATAT by $1.3\%$.
Conclusions. SwinV2, a pre-trained VT model, effectively classifies photometric light curves, outperforming traditional models and offering a promising approach for large-scale surveys. This highlights the potential of using visual representations of light curves, with future prospects including the integration of tabular data, textual information, and multi-modal learning to enhance analysis and classification in time-domain astronomy.
@@ 1 Introduction
The Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al. 2019) is set to revolutionize our understanding of the universe by capturing an unprecedented volume of data, with an average of 10 million alerts and $\sim 20\,\text{TB}$ of raw data every night (Hambleton et al. 2023). This vast dataset, covering a range of celestial phenomena, including transient events and stellar variability, will provide astronomers with a unique opportunity to study the physical properties and evolution of a diverse array of astronomical objects. However, the scale of these observations necessitates the development of efficient and accurate classification tools, as spectroscopic follow-up will be limited. Automatic classification methods, particularly those based on photometric data, will be essential to distinguish between different types of variable stars and transient phenomena, ensuring that the scientific potential of the LSST data is fully realized (Fraga et al. 2024).
@@
Machine learning and deep learning techniques have played a central role in advancing the classification of light curves (Pichara et al. 2012; Kim et al. 2011; Villar et al. 2019; Becker et al. 2020; Cabrera-Vives et al. 2024). Traditional machine learning approaches have been highly effective in classifying variable stars and transient events by relying on hand-engineered features extracted from light curves. For example, Richards et al. (2011) and Kim et al. (2014) employed statistical and time-domain features to enhance classification accuracy. Karpenka et al. (2013) combined parametric functional fitting of supernova light curves with a machine learning algorithm, yielding accurate and robust classifications. Lochner et al. (2016) developed a multi-stage pipeline that evaluates feature extraction techniques and machine learning models. Boone (2019) improved classification by training a boosted decision tree on features extracted from the augmented light curves, achieving state-of-the-art performance in the LSST Astronomical Time-Series Classification Challenge (PLAsTiCC; Kessler et al. 2019). However, this feature engineering process is resource-intensive and requires domain knowledge to be applied (Donalek et al. 2013; Graham et al. 2014; Nun et al. 2017; Sánchez-Sáez et al. 2021). Deep learning addresses this limitation by automatically learning relevant representations directly from light curves, removing the need for manual feature extraction. This advantage has led to the widespread adoption of recurrent neural networks (RNNs; Rumelhart et al. 1986), long short-term memory networks (LSTMs; Hochreiter & Schmidhuber 1997), and gated recurrent units (GRUs; Cho et al. 2014) (Charnock & Moss 2017; Naul et al. 2018; Muthukrishna et al. 2019; Carrasco-Davis et al. 2019; Möller & de Boissière 2020; Becker et al. 2020; Jamal & Bloom 2020; Gómez et al. 2020; Donoso-Oliva et al. 2021), as well as transformer-based models, which have shown great promise in handling astronomical time series data (Morvan et al. 2022; Pan et al. 2022; Pimentel et al. 2022; Donoso-Oliva et al. 2023; Moreno-Cartagena et al. 2023; Cádiz-Leyton et al. 2024, 2025; Allam Jr & McEwen 2024; Cabrera-Vives et al. 2024), making them increasingly popular for astronomical classification tasks.
@@
Nevertheless, even with deep learning models in the time domain, careful preprocessing is required when dealing with multi-band observations and irregular time sampling, as each band is often observed at different, non-uniform time intervals. Various techniques have been developed to address these challenges, such as uniform sampling imputation between previous and subsequent observations (Charnock & Moss 2017), linear interpolation on a regular time grid across all bands (Muthukrishna et al. 2019), Gaussian Process (GP) interpolation (Boone 2019; Villar et al. 2020; Jamal & Bloom 2020), masking mechanisms to indicate the availability of each band during observations (Möller & de Boissière 2020; Pimentel et al. 2022; Cabrera-Vives et al. 2024), and ensembles of RNNs that integrate information from different bands into a single representation (Becker, I. et al. 2025). While these methods have been evaluated on various datasets, their generalization across different surveys and observational conditions remains an open question (Moreno-Cartagena et al. 2023). Additionally, many of these approaches require training from scratch and hyperparameter tuning to search for the optimal architecture for a given task.
@@
Foundational models, which are trained only once on massive datasets in a self-supervised manner (pre-training) to learn general domain representations and later adapted for specific tasks (fine-tuning), have demonstrated exceptional generalization across diverse downstream applications in natural language processing and computer vision, such as text generation (Brown et al. 2020), machine translation (Vaswani et al. 2017), image classification (Dosovitskiy et al. 2021), and object detection (Carion et al. 2020), among others. Notable examples include GPT (Radford et al. 2018) and BERT (Devlin et al. 2019) for text, as well as Vision Transformer (ViT; Dosovitskiy et al. 2021) for images, each specializing in a single modality. In the context of astronomical time series data, Astromer (Donoso-Oliva et al. 2023) represents an early effort to pre-train a transformer on large-scale real-world light curve data, positioning it as a potential foundational model for time-domain astronomy. However, such pre-training requires significant computational resources, optimized architectural design, and carefully curated datasets, often taking several days to converge.
@@
Given these constraints, we propose leveraging the knowledge embedded in pre-trained visual transformer (VT) models for the analysis and classification of multi-band photometric light curves. These models have demonstrated the ability to capture fine details in images across various tasks (Han et al. 2023; Khan et al. 2022). By using the informative representations learned from extensive pre-training on large-scale datasets, conducted on powerful GPU clusters, we can establish a starting point for fine-tuning the model on light curve images, bypassing the need to use significant computational resources for pre-training and searching for an optimal architecture. Moreover, these models can directly process raw multi-band light curves as images, offering a straightforward yet highly effective approach to handling the data while avoiding complex feature engineering or multi-band preprocessing.
@@
A few studies considered light curves in the time domain for visual models. Mahabal et al. (2017) transformed light curves from the Catalina Real-time Transient Survey (CRTS; Drake et al. 2012) into 2D images by creating magnitude differences and time differences between pairs of points as input for a convolutional neural network (CNN; LeCun et al. 1998). Pasquet-Itam & Pasquet (2018) and Pasquet et al. (2019) used CNNs for the classification of astronomical events, representing light curves as images, where the vertical axis corresponds to the number of photometric bands, and the horizontal axis, represents the number of days in the first study and the number of observations in the second. Qu et al. (2021) and Qu & Sako (2022) transformed multi-band light curves into 2D heatmaps using GPs to represent flux and its uncertainties over time and wavelength. These heatmaps were stacked into $32{\times}180{\times}2$ tensors and fed into a CNN for supernova classification. Szklenár et al. (2020, 2022) explored the classification of phase-folded light curves using the Optical Gravitational Lensing Experiment (OGLE)-III and -IV databases (Udalski et al. 2008, 2015), representing them as $128{\times}128$ monochromatic pixel images with brightness measurements plotted as white dots on a black background. Similarly, Monsalves et al. (2024) transformed light curves into $32{\times}32$ pixel 2D histograms, where phase and magnitude bins were used as inputs to a CNN to classify different types of variable stars. However, all these studies depended on a prior level of preprocessing and did not use pre-trained VT models.
@@
In this work, we present the use of a pre-trained VT model called Swin Transformer V2 (SwinV2; Liu et al. 2022), which we evaluate on two datasets: the public Catalog of Variable Stars (Alcock et al. 2003), derived from the Massive Compact Halo Object (MACHO; Alcock et al. 2000) survey, and the first round of the recent Extended LSST Astronomical Time-Series Classification Challenge (ELAsTiCC [^1] The DESC ELAsTiCC Challenge.; Narayan & Team 2023), a multi-band dataset of photometric light curves for variable stars and supernovae designed to simulate the observations of the future LSST. Our work builds upon the methodology proposed by Li et al. (2024), which focuses on benchmark datasets for time series. We applied this method to light curves, performed hyperparameter tuning, and evaluated its performance using both single- and multi-band configurations for the MACHO dataset and six bands for the ELAsTiCC dataset. Additionally, we assess its performance in scenarios with a limited number of astronomical objects per class and imbalanced datasets. To our knowledge, this is the first study to apply a pre-trained VT to raw multi-band light curves in the time domain.
@@
This paper is organized as follows: In Section 2, we describe the methodology for converting multi-band light curves into images and present the architecture of the SwinV2 visual transformer model. Section 3 details the datasets and experimental setup. In Section 4, we present the results of our classification experiments on the MACHO and ELAsTiCC datasets, including comparisons with established methods. Finally, Section 5 summarizes the main conclusions of this work and suggests directions for future research.
@@ 2.1 Multi-band light curves to images
Transforming time series into images is crucial for applying VT models, which have proven effective in analyzing complex visual data (Dosovitskiy et al. 2021; Liu et al. 2021, 2023). In this study, we evaluate two approaches for converting multi-band light curves into images: Grid and Overlay. Figure 1 (a) shows the input for the Grid approach, while (b) illustrates the input for the Overlay approach.
@@
In the Grid approach, light curves from different bands are arranged in a grid, with each band displayed in a separate panel, following a similar methodology to Li et al. (2024). We use a square layout that adjusts to the number of bands present in each light curve, leaving empty panels when there are no observations for a particular band. This arrangement facilitates a clear comparison between bands by keeping them visually distinct. In the Overlay approach, all bands of a light curve are plotted together within the same space. This approach allows us to examine how interactions among observations from different bands influence the visual representation and model performance.
@@
Figure 1: Comparison of visualization strategies. Panel (a) illustrates the Grid approach, while panel (b) depicts the Overlay approach. Each image was generated using the best hyperparameters identified for its respective method. The single-band and two-band example is a Long-Period Variable (LPV) (ID: 24.3466.13) from the MACHO dataset, whereas the six-band example is a Pair-Instability Supernova (PISN) (ID: 41820707) from the ELAsTiCC dataset.
@@
To ensure that all light curve images are on the same scale and maintain consistent x- and y-limits across different images, we apply Min-Max normalization to each entire light curve (considering all bands). This approach preserves the relative differences among bands within each light curve while scaling time and flux values across all light curves to the range [0, 11]. As a consequence, the global temporal and brightness context across light curves is removed, which encourages the model to focus on light curve shape to discriminate between astronomical objects. Global normalization was avoided, as it suppressed low-amplitude variability and negatively affected classification performance. Axis labels and ticks are also removed before saving the plots. Additionally, the flux errors are scaled using the same Min-Max factor applied to the flux, maintaining the relative uncertainty information. This method ensures that all images are visually comparable, with uniform axis limits and scale throughout the dataset. Visual examples of these representation approaches applied to randomly selected light curves from the test set are provided in Appendix A, in Fig. 10 and Fig. 11, offering a broader view of the differences between the One-band, Overlay, and Grid methods across both the MACHO and ELAsTiCC datasets.
@@
Since pre-trained VT models are trained on RGB images, the input must conform to this expected format, requiring a three-channel representation. In the Grid approach, the specific color assigned to each band is inconsequential, as each band occupies a distinct spatial position. However, in the Overlay approach, where all bands are combined into the same visual space, color selection becomes crucial to avoid introducing visual biases that could affect the model’s performance. To mitigate this issue and ensure that all bands contribute equally to the final representation, we assign colors that approximate distinct RGB components, maintaining a perceptually balanced distribution across the spectrum. For datasets with six bands, we allocate two bands to the red spectrum with the RGB values (255, 0, 127) and (255, 127, 0), two bands to the green spectrum with (0, 255, 127) and (127, 255, 0), and the remaining two to the blue spectrum with (0, 127, 255) and (127, 0, 255). This selection ensures an even distribution of color intensity, preventing any single band from dominating the representation. For datasets containing only two bands, each is mapped to a primary RGB component to preserve visual balance and maintain interpretability.
@@
2.2 Visual transformer models
VT models represent a significant evolution in deep learning architectures for image processing. Unlike conventional approaches based on convolutions, such as CNNs, VTs employ self-attention mechanisms to capture long-range relationships between different regions of an image (Dosovitskiy et al. 2021). This methodology enables VTs to capture intricate structures within images, achieving robust performance in various tasks, such as classification (Liu et al. 2021), object detection (Carion et al. 2020), semantic and instance segmentation (Wang et al. 2021; Cheng et al. 2021), object tracking (Chen et al. 2021b), image generation (Jiang et al. 2021), and image enhancement (Chen et al. 2021a), among others. Furthermore, VT models that are pre-trained can use the prior knowledge gained from extensive datasets, making them adaptable to particular tasks via a fine-tuning process.

### v4

# Leveraging pre-trained vision Transformers for multi-band photometric light curve classification

@@ Abstract
Aims. This study investigates the potential of a pre-trained vision Transformer (VT) model, specifically the Swin Transformer V2 (SwinV2), to classify photometric light curves without the need for feature extraction or multi-band preprocessing. The goal is to assess whether this image-based approach can accurately differentiate astronomical phenomena and if it can serve as a viable option for working with multi-band photometric light curves.
Methods. We transformed each multi-band light curve into an image. These images served as input to the SwinV2 model, which was pre-trained on ImageNet-21K. The datasets employed include the public Catalog of Variable Stars from the Massive Compact Halo Object (MACHO) survey, using both one and two bands, and the first round of the recent Extended LSST Astronomical Time-Series Classification Challenge (ELAsTiCC), which includes six bands. The model’s performance was evaluated based on six classes for the MACHO dataset and 20 distinct classes of variable stars and transient events for the ELAsTiCC dataset.
Results. The fine-tuned SwinV2 model achieved a better performance than models specifically designed for light curves, such as Astromer and the Astronomical Transformer for time series And Tabular data (ATAT). When trained on the “full dataset” of MACHO, it attained a macro F1-score of 80.2% and outperformed Astromer in single-band experiments. Incorporating a second band further improved performance, increasing the F1-score to $84.1\%$. In the ELAsTiCC dataset, SwinV2 achieved a macro F1-score of $65.5\%$, slightly surpassing ATAT by $1.3\%$.
Conclusions. SwinV2, a pre-trained VT model, effectively classifies photometric light curves. It outperforms traditional models and offers a promising approach for large-scale surveys. This highlights the potential of using visual representations of light curves, with future prospects including the integration of tabular data, textual information, and multi-modal learning to enhance analysis and classification in time-domain astronomy.
@@ 1 Introduction
The Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al. 2019) is set to revolutionize our understanding of the Universe by capturing an unprecedented volume of data, with an average of 10 million alerts and $\sim 20\,\text{TB}$ of raw data every night (Hambleton et al. 2023). This vast dataset covers a range of celestial phenomena, including transient events and stellar variability, and will provide astronomers with a unique opportunity to study the physical properties and evolution of a diverse array of astronomical objects. However, the scale of these observations necessitates the development of efficient and accurate classification tools, as spectroscopic follow-up will be limited. Automatic classification methods, particularly those based on photometric data, will be essential for distinguishing between different types of variable stars and transient phenomena, ensuring that the scientific potential of the LSST data is fully realized (Fraga et al. 2024).
@@
Machine learning and deep learning techniques have played a central role in advancing the classification of light curves (Pichara et al. 2012; Kim et al. 2011; Villar et al. 2019; Becker et al. 2020; Cabrera-Vives et al. 2024). Traditional machine learning approaches have been highly effective in classifying variable stars and transient events by relying on hand-engineered features extracted from light curves. For example, Richards et al. (2011) and Kim et al. (2014) employed statistical and time-domain features to enhance classification accuracy. Karpenka et al. (2013) combined parametric functional fitting of supernova light curves with a machine learning algorithm, yielding accurate and robust classifications. Lochner et al. (2016) developed a multi-stage pipeline that evaluates feature extraction techniques and machine learning models. Boone (2019) improved classification by training a boosted decision tree on features extracted from the augmented light curves, achieving state-of-the-art performance in the LSST Astronomical Time-Series Classification Challenge (PLAsTiCC; Kessler et al. 2019). However, this feature engineering process is resource-intensive and requires domain knowledge to be applied (Donalek et al. 2013; Graham et al. 2014; Nun et al. 2017; Sánchez-Sáez et al. 2021). Deep learning addresses these limitations by automatically learning relevant representations directly from light curves, removing the need for manual feature extraction. This advantage has led to the widespread adoption of recurrent neural networks (RNNs; Rumelhart et al. 1986), long short-term memory networks (LSTMs; Hochreiter & Schmidhuber 1997), and gated recurrent units (GRUs; Cho et al. 2014) (Charnock & Moss 2017; Naul et al. 2018; Muthukrishna et al. 2019; Carrasco-Davis et al. 2019; Möller & de Boissière 2020; Becker et al. 2020; Jamal & Bloom 2020; Gómez et al. 2020; Donoso-Oliva et al. 2021) as well as transformer-based models, which have shown great promise in handling astronomical time series data (Morvan et al. 2022; Pan et al. 2022; Pimentel et al. 2022; Donoso-Oliva et al. 2023; Moreno-Cartagena et al. 2023; Cádiz-Leyton et al. 2024, 2025; Allam Jr & McEwen 2024; Cabrera-Vives et al. 2024), making them increasingly popular for astronomical classification tasks.
@@
Nevertheless, even with deep learning models in the time domain, careful preprocessing is required when dealing with multi-band observations and irregular time sampling, as each band is often observed at different non-uniform time intervals. Various techniques have been developed to address these challenges, such as uniform sampling imputation between previous and subsequent observations (Charnock & Moss 2017), linear interpolation on a regular time grid across all bands (Muthukrishna et al. 2019), Gaussian process (GP) interpolation (Boone 2019; Villar et al. 2020; Jamal & Bloom 2020), masking mechanisms to indicate the availability of each band during observations (Möller & de Boissière 2020; Pimentel et al. 2022; Cabrera-Vives et al. 2024), and ensembles of RNNs that integrate information from different bands into a single representation (Becker, I. et al. 2025). While these methods have been evaluated on various datasets, their generalization across different surveys and observational conditions remains an open question (Moreno-Cartagena et al. 2023). Additionally, many of these approaches require training from scratch and hyperparameter tuning to search for the optimal architecture for a given task.
@@
Foundational models, which are trained only once on massive datasets in a self-supervised manner (pre-training) to learn general domain representations and later adapted for specific tasks (fine-tuning), have demonstrated exceptional generalization across diverse downstream applications in natural language processing and computer vision, such as text generation (Brown et al. 2020), machine translation (Vaswani et al. 2017), image classification (Dosovitskiy et al. 2021), and object detection (Carion et al. 2020), among others. Notable examples include GPT (Radford et al. 2018) and BERT (Devlin et al. 2019) for text as well as Vision Transformer (ViT; Dosovitskiy et al. 2021) for images, with each model specializing in a single modality. In the context of astronomical time series data, Astromer (Donoso-Oliva et al. 2023) represents an early effort to pre-train a transformer on large-scale real-world light curve data, positioning it as a potential foundational model for time-domain astronomy. However, such pre-training requires significant computational resources, optimized architectural design, and carefully curated datasets, often taking several days to converge.
@@
Given these constraints, we propose leveraging the knowledge embedded in pre-trained vision Transformer (VT) models for the analysis and classification of multi-band photometric light curves. These models have demonstrated the ability to capture fine details in images across various tasks (Han et al. 2023; Khan et al. 2022). By using the informative representations learned from extensive pre-training on large-scale datasets conducted on powerful GPU clusters, we can establish a starting point for fine-tuning the model on light curve images, bypassing the need to use significant computational resources for pre-training and searching for an optimal architecture. Moreover, these models can directly process raw multi-band light curves as images, offering a straightforward yet highly effective approach to handling the data while avoiding complex feature engineering or multi-band preprocessing.
@@
A few studies have considered light curves in the time domain for visual models. Mahabal et al. (2017) transformed light curves from the Catalina Real-time Transient Survey (CRTS; Drake et al. 2012) into 2D images by creating magnitude differences and time differences between pairs of points as input for a convolutional neural network (CNN; LeCun et al. 1998). Pasquet-Itam & Pasquet (2018) and Pasquet et al. (2019) used CNNs for the classification of astronomical events, representing light curves as images, where the vertical axis corresponds to the number of photometric bands and the horizontal axis represents the number of days in the first study and the number of observations in the second. Qu et al. (2021) and Qu & Sako (2022) transformed multi-band light curves into 2D heatmaps using GPs to represent flux and its uncertainties over time and wavelength. These heatmaps were stacked into $32{\times}180{\times}2$ tensors and fed into a CNN for supernova classification. Szklenár et al. (2020, 2022) explored the classification of phase-folded light curves using the Optical Gravitational Lensing Experiment (OGLE)-III and -IV databases (Udalski et al. 2008, 2015), representing them as $128{\times}128$ monochromatic pixel images with brightness measurements plotted as white dots on a black background. Similarly, Monsalves et al. (2024) transformed light curves into $32{\times}32$ pixel 2D histograms, where phase and magnitude bins were used as inputs to a CNN to classify different types of variable stars. However, all of these studies depended on a prior level of preprocessing and did not use pre-trained VT models.
@@
In this work, we present the use of a pre-trained VT model called Swin Transformer V2 (SwinV2; Liu et al. 2022), which we evaluate on two datasets: the public Catalog of Variable Stars (Alcock et al. 2003), derived from the Massive Compact Halo Object (MACHO; Alcock et al. 2000) survey, and the first round of the recent Extended LSST Astronomical Time-Series Classification Challenge (ELAsTiCC [^1] The DESC ELAsTiCC Challenge.; Narayan & Team 2023), a multi-band dataset of photometric light curves for variable stars and supernovae designed to simulate the observations of the future LSST. Our work builds on the methodology proposed by Li et al. (2024), which focuses on benchmark datasets for time series. We applied this method to light curves, performed hyperparameter tuning, and evaluated its performance using both single- and multi-band configurations for the MACHO dataset and six bands for the ELAsTiCC dataset. Additionally, we assessed its performance in scenarios with a limited number of astronomical objects per class and imbalanced datasets. To our knowledge, this is the first study to apply a pre-trained VT to raw multi-band light curves in the time domain.
@@
This paper is organized as follows: In Section 2, we describe the methodology for converting multi-band light curves into images and present the architecture of the SwinV2 vision Transformer model. Section 3 provides details of the datasets and experimental setup. In Section 4, we present the results of our classification experiments on the MACHO and ELAsTiCC datasets, including comparisons with established methods. Finally, in Section 5 we summarize the main conclusions of this work and suggest directions for future research.
@@ 2.1 Multi-band light curves to images
Transforming time series into images is crucial for applying VT models, which have proven effective in analyzing complex visual data (Dosovitskiy et al. 2021; Liu et al. 2021, 2023). In this study, we evaluate two approaches for converting multi-band light curves into images: “Grid” and “Overlay”. Figure 1 (a) shows the input for the “Grid” approach, while (b) illustrates the input for the “Overlay” approach.
@@
In the “Grid” approach, light curves from different bands are arranged in a grid, with each band displayed in a separate panel, following a similar methodology to Li et al. (2024). We use a square layout that adjusts to the number of bands present in each light curve, leaving empty panels when there are no observations for a particular band. This arrangement facilitates a clear comparison between bands by keeping them visually distinct. In the “Overlay” approach, all bands of a light curve are plotted together within the same space. This approach allows us to examine how interactions among observations from different bands influence the visual representation and model performance.
@@
Figure 1: Comparison of visualization strategies. Panel (a) illustrates the “Grid” approach, while panel (b) depicts the “Overlay” approach. Each image was generated using the best hyperparameters identified for its respective method. The single-band and two-band example is a Long-Period Variable (LPV) (ID: 24.3466.13) from the MACHO dataset, whereas the six-band example is a Pair-Instability Supernova (PISN) (ID: 41820707) from the ELAsTiCC dataset.
@@
To ensure that all light curve images are on the same scale and maintain consistent x- and y-limits across different images, we applied min-max normalization to each entire light curve (considering all bands). This approach preserves the relative differences among bands within each light curve while scaling time and flux values across all light curves to the range [0, 11]. As a consequence, the global temporal and brightness context across light curves is removed, which encourages the model to focus on light curve shape to discriminate between astronomical objects. Global normalization was avoided, as it suppressed low-amplitude variability and negatively affected classification performance. Axis labels and ticks are also removed before saving the plots. Additionally, the flux errors are scaled using the same Min-Max factor applied to the flux, maintaining the relative uncertainty information. This method ensures that all images are visually comparable, with uniform axis limits and scale throughout the dataset. Visual examples of these representation approaches applied to randomly selected light curves from the test set are provided in Appendix A, in Fig. 10 and Fig. 11, offering a broader view of the differences between the One-band, “Overlay”, and “Grid” methods across both the MACHO and ELAsTiCC datasets.
@@
Since pre-trained VT models are trained on RGB images, the input must conform to this expected format, requiring a three-channel representation. In the “Grid” approach, the specific color assigned to each band is inconsequential, as each band occupies a distinct spatial position. However, in the “Overlay” approach, where all bands are combined into the same visual space, color selection becomes crucial to avoid introducing visual biases that could affect the model’s performance. To mitigate this issue and ensure that all bands contribute equally to the final representation, we assign colors that approximate distinct RGB components, maintaining a perceptually balanced distribution across the spectrum. For datasets with six bands, we allocate two bands to the red spectrum with the RGB values (255, 0, 127) and (255, 127, 0), two bands to the green spectrum with (0, 255, 127) and (127, 255, 0), and the remaining two to the blue spectrum with (0, 127, 255) and (127, 0, 255). This selection ensures an even distribution of color intensity, preventing any single band from dominating the representation. For datasets containing only two bands, each is mapped to a primary RGB component to preserve visual balance and maintain interpretability.
@@
2.2 Vision Transformer models
Vision Transformer models represent a significant evolution in deep learning architectures for image processing. Unlike conventional approaches based on convolutions, such as CNNs, VTs employ self-attention mechanisms to capture long-range relationships between different regions of an image (Dosovitskiy et al. 2021). This methodology enables VTs to capture intricate structures within images, achieving robust performance in various tasks, such as classification (Liu et al. 2021), object detection (Carion et al. 2020), semantic and instance segmentation (Wang et al. 2021; Cheng et al. 2021), object tracking (Chen et al. 2021b), image generation (Jiang et al. 2021), and image enhancement (Chen et al. 2021a), among others. Furthermore, VT models that are pre-trained can use the prior knowledge gained from extensive datasets, making them adaptable to particular tasks via a fine-tuning process.